# Replacing Softmax Similarity with a Sharpened Angular Similarity: Theory and Practice of Scaling to Billion-Context Attention

### v2

@@ 4 Experiments
To avoid cherry-picking and ensure comparability, we adopt the evaluation suites standard in prior efficient-attention work—Linear Attention, Linformer, and Random Feature Attention (RFA) [15, 32, 24]. Concretely, we include text classification (QNLI [27], SST-2 [28], IMDB [20], and Yahoo [37]) to probe moderate-length discriminative accuracy (as in Linformer); autoregressive language modeling to test token-level modeling (as in RFA) on WikiText-103 [22] and Penn Tree Bank (PTB) corpus [21]; masked language modeling on Tiny Stories dataset [13]; image classification (CIFAR-10 [17] and FashionMNIST [33]) to test expressivity using Vision Transformer [12] architecture (as in Linear Attention); and long-context reasoning via Long Range Arena benchmark [29] (e.g., ListOps and Text Retrieval) to stress scaling and accuracy. Together these cover four regimes—bidirectional, autoregressive, long-context, and moderate-context text/image classification. Beyond these standard benchmarks, we additionally report extreme-length scaling experiments to tens of millions of tokens. To the best of our knowledge, this is the first work to report experiments with attention spanning close to 100 million tokens. In this section, we introduce an additional hyperparameter $M$ representing the number of ensembles per head i.e., independent replications of the whole sketching scheme inside each attention head. For clarity, FlashAttention is an exact, fused-kernel implementation of Softmax Attention; we therefore use “FlashAttention” and “Softmax Attention” interchangeably when discussing accuracy and runtime.
(a) RACE - $P=2,L=2,M=1$
(b) RACE - $P=3,L=3,M=1$
(c) RACE - $P=4,L=4,M=1$
(d) RACE - $P=2,L=2,M=1$
(e) RACE - $P=3,L=3,M=1$
(f) RACE - $P=4,L=4,M=1$
Figure 3: A rigorous scaling stress-test across hardware. The top row shows GPU scaling results; the bottom row shows CPU scaling results. We run a single forward-backward pass configured with 1 batch, 4 heads, and an embedding dimension of 128. Linformer and Performer use the same low-rank/feature dimension as in Table 1.
Table 1: IMDB @ N=512 results.
Method P L M Accuracy
RACE 1 2 1 81.3%
RACE 2 2 2 80.6%
RACE 3 3 1 81.3%
RACE 4 4 1 81.3%
Linformer-128 78.2%
FlashAttention 80.0%
Angular ($\gamma$=6) 79.6%
Linear 80.9%
Performer-256 81%
Table 2: Tiny Stories (subset) @ N=512
Method P L M Perplexity
RACE 3 4 1 3.9
RACE 4 4 1 3.3
RACE 5 4 2 2.7
RACE 5 5 1 5.1
Linear 6
Angular ($\gamma$ = 8) 2.9
FlashAttention 3.1
Linformer-128 4.6
Performer-256 7.1
Table 3: ListOps @ N=2000
Method P L M Acc.
RACE 2 2 1 41.9%
RACE 2 3 2 41.0%
RACE 3 3 1 41.3%
RACE 4 3 1 41.2%
Linformer-128 38.9%
FlashAttention 41.4%
Angular ($\gamma$=8) 42.2%
Linear 39.6%
Performer-256 40.2%
Table 4: QNLI @ N=2048
Method P L M Accuracy
RACE 2 2 1 60.7%
RACE 3 3 1 60.7%
RACE 4 4 1 61.1%
RACE 5 5 1 60.4%
Linformer-128 60.6%
Linear 60.7%
FlashAttention 61.1%
Angular ($\gamma$=8) 61.7%
Performer-256 61.0%
Table 5: Text Retrieval @ N=8000
Method P L M Acc.
RACE 2 2 1 80.3%
RACE 2 3 1 80.5%
RACE 3 3 1 80.8%
RACE 4 4 1 80.9%
Linformer-128 76.1%
Linear 80.6%
Performer-256 80.8%
Table 6: SST-2 @ N=1024
Method P L M Accuracy
RACE 2 2 1 76.7%
RACE 4 4 1 79.4%
Linformer-128 75.1%
Linear 78 %
FlashAttention 78.5%
Angular ($\gamma$=8) 77.2%
Performer-256 77.3%
Table 7: PTB @ N=128
Method P L M Test PPL
RACE 2 2 1 54.7
RACE 3 3 1 54.2
RACE 4 4 1 53.4
Angular ($\gamma$=8) 58.8
Angular ($\gamma$=12) 57.6
Linear 73.2
FlashAttention 55.4
Table 8: WikiText-103 @ N=1024
Method P L M Test PPL
RACE 2 2 1 23.9
RACE 2 3 1 23.4
RACE 3 3 1 21.9
RACE 3 4 1 21.5
RACE 4 4 1 20.9
FlashAttention 20.9
Angular ($\gamma$=8) 19
Table 9: Sequential CIFAR-10 @ N=1024
Method P L M Accuracy
RACE 2 2 1 63.7%
RACE 3 3 1 62.5%
RACE 3 5 1 65.7%
RACE 4 5 1 65.9%
Linformer-128 63.7%
FlashAttention 61.44%
Angular ($\gamma$=8) 61.69%
Linear 60%
Performer-256 64.9%
Table 10: Sequential FashionMNIST @ N=784
Method P L M Accuracy
RACE 2 5 1 87.7%
RACE 3 5 1 87.5%
RACE 4 4 1 86.6%
RACE 4 5 1 85.7%
Linformer-128 87.7%
FlashAttention 87.2%
Angular ($\gamma$=8) 86.4%
Linear 85.8%
Performer-256 86.6%
Table 11: Tiny Stories @ N=1024
Method P L M Perplexity
RACE 2 2 1 4.2
RACE 2 3 1 4
RACE 3 3 1 3.2
RACE 4 4 1 2.6
Linear 7
Linformer-128 3.7
FlashAttention 2.7
Angular ($\gamma$ = 8) 2.5
FAVOR+ 10
Table 12: Yahoo @ N=256
Method P L M Accuracy
RACE 2 2 1 66.9%
RACE 3 3 1 66.6%
RACE 4 3 1 66.6%
RACE 4 4 1 67.2%
Linformer-128 64.7%
FlashAttention 67.2%
Angular ($\gamma$=8) 67.0%
Linear 66.9%
Performer-256 64.9%
Baselines: We compare RACE against widely used baselines with publicly available implementations: FlashAttention [10], Linear Attention [15], Performer [6], and Linformer [32]. These span exact, kernel-linear, and low-rank approximations. All models are tuned per authors’ guidelines and trained under identical settings.
4.1 Is RACE Attention as Accurate as Transformers?
We report text-classification accuracies in Tables 1, 4, 6, and 12; long-context (LRA) results—ListOps and Text Retrieval—in Tables 3 and 5; image-classification accuracies in Table 9 and 10; and masked language modeling (MLM) perplexities in Tables 2 and 11. For autoregressive language modeling, RACE matches softmax-level perplexity on WikiText-103 and improves upon it on PTB (Tables 7, 8). These results indicate that RACE preserves accuracy in the overlapping regime while delivering consistent gains on moderate-context and long-context settings.
Unless stated otherwise, all methods use the same Transformer backbone (layers, heads, embedding dimension, dropout) and training budget. We train with identical optimizers, schedulers, and batch sizes; full hyperparameters appear in Table 13. Metrics are reported from the best-validation checkpoint. Despite the extra parameters introduced by Linformer’s lengthwise projections, it does not outperform RACE under matching training conditions. For the long-range task in Table 5, FlashAttention and Angular Attention have quadratic time complexity and are prohibitively slow; we therefore exclude them. All experiments were run on NVIDIA A100 GPUs.
4.2 Can we reach 100 million context window on popular hardware?
In this section, we evaluate how RACE Attention scales across common hardware relative to strong baselines. For RACE, we use sketch parameters $(P,L,M)$ chosen to match FlashAttention’s accuracy/perplexity on the same tasks in Section 4.1. For each method, we measure the wall-clock time for a single forward–backward pass of the multi-head attention layer with 1 batch, 4 heads, and embedding dimension of 128, as a function of sequence length, stress-testing context lengths up to 100 million tokens.
(a) GPU scaling
(b) CPU scaling
(c) $d$ scaling @ N=65536N{=}65536
Figure 4: A rigorous scaling stress-test (including FlashAttention) across hardware. Plots (a)–(b) use logarithmic axes. RACE is evaluated with (P=2,L=2,M=1)(P{=}2,L{=}2,M{=}1) throughout; Linformer and Performer use the same low-rank/feature dimension as in Table 1.
How far can we scale attention on standard Intel Xeon® Gold 5220R CPU?
RACE scales to 75 million tokens for a single forward-backward pass on CPU. By contrast, FlashAttention becomes prohibitively slow at $\sim\!2$ million tokens due to the quadratic time scaling in sequence length $N$ (see figs. 3 and 4). It is worth noting that FlashAttention does not run out of memory on the CPU DRAM. RACE is more than $10000\times$ faster than FlashAttention at context length of $\sim\!33$ million. RACE finishes comfortably under 10 seconds for a single forward-backward pass on this hardware while FlashAttention takes approximately $10^{5}$ seconds on the same hardware. Although the absolute runtime grows with $N$, RACE’s speedup over FlashAttention increases. At 75 million tokens, RACE finishes in about 100 seconds. This is expected because RACE is linear and FlashAttention is quadratic in $N$. The experiments also highlight that linear attentions’ approximations are not only inaccurate but also significantly slower and have large memory overheads due to large hidden constants. They run about an order of magnitude slower than RACE Attention and even go out of memory at $\sim\!33$ million tokens.
How far can we scale attention on the most powerful GH200 GPU?
An NVIDIA GH200 has 96GB of memory. Here, we observe a similar trend. RACE scales up to 12 million tokens for a single forward-backward pass, whereas FlashAttention becomes impractical around $\sim\!4$ million tokens (see figs. 3 and 4). At $\sim\!4$ million tokens RACE takes merely 0.1 seconds to finish, while FlashAttention needs about 500 seconds, making RACE about $5000\times$ faster on GPUs. While FlashAttention’s activation memory scales linearly with $N$ and $d$, the GPU’s high-bandwidth memory (HBM) is nevertheless exhausted for sufficiently large $N$. This is because, we must retain $Q,K,V,O$ (and their gradients) of size $\mathcal{O}(BHNd)$ with large constants. Even though the $N\times N$ score matrix is never materialized, this footprint exceeds HBM capacity at large $N$, leading to out-of-memory failures (see fig. 4). Furthermore, RACE even scales better on GPU than the cheap but less accurate linear baselines, and they run out of memory around $\sim\!4$ million tokens. RACE handles about $3.5\times$ longer contexts than FlashAttention.
4.3 Right Algorithm beats Hardware Acceleration!
Figure 5: A rigorous scaling test for algorithmic comparison between FlashAttention on GPU vs. RACE Attention on CPU
[…]
@@ 5 Conclusion
We introduced RACE Attention, a linear-time, memory-efficient alternative to Softmax Attention that estimates a sharpened angular kernel via RACE sketches. By replacing explicit pairwise scores with aggregated bucket scores, compute and activation memory scale linearly with context length $N$, and embedding dimension $d$, with constants governed by the sketch parameters $(P,L,M)$. In our rigorous stress-test, we scale at context lengths up to $\mathbf{75\,million}$ tokens on CPU and $\mathbf{12\,million}$ tokens on GPU (single forward–backward pass), a regime infeasible for Softmax Attention and other highly optimized linear attention mechanisms under comparable settings.
At moderate $N$, optimized GPU Softmax Attention (FlashAttention) yields lower wall-clock time due to high parallelism and effective kernel fusion. As $N$ grows, however, quadratic cost and activation memory dominate, capping sequence length and increasing runtime. In contrast, CPU-based RACE maintains linear scaling, enabling both longer contexts and faster speed in the long-context regime by aggregating values within hash buckets. It is promising to extend RACE Attention for inference only scenarios, potentially eliminating the need for a K cache, and to develop a CUDA kernel, analogous to our OpenMP implementation, that scales well for autoregressive tasks on GPU. We leave these explorations for future work.
@@ Acknowledgments
The work was primarily supported by Rice Ken Kennedy Institute (K2I) Generative AI Cluster Funding. We are grateful to Dr. Aditya Desai and Dr. Zhaozhuo Xu for their valuable insights and thought-provoking discussions.
@@
Unless otherwise stated, we use a linear warmup–decay learning-rate schedule. Let $T$ denote the total number of optimizer updates ($T=\mathrm{epochs}\times\mathrm{len(train\_loader)}$)). The learning rate increases linearly from 0 to the base value (provided in Table 13) over the first $0.01T$ updates, then decays linearly to 0 over the remaining T0.01TT-0.01T updates; the scheduler is stepped once per optimizer update. We use the AdamW optimizer with hyperparameters listed in Table 13. For learning rate $2e^{-4}$, we don’t use linear warmup or a scheduler, we simply train the model with the constant learning rate for certain number of epochs given in Table 13.

### v3

@@ 4 Experiments
To ensure comparability and avoid cherry-picking, we adopt the standard evaluation suites used in prior efficient-attention work—Linear Attention, Linformer, and Random Feature Attention (RFA) [18, 38, 27]. We evaluate text classification on QNLI, SST-2, IMDB, Yahoo, and Arxiv [31, 34, 23, 44, 17]; autoregressive language modeling on WikiText-103 and PTB [25, 24]; masked language modeling on Tiny Stories [15]; image classification on CIFAR-10, FashionMNIST, and Food-101 using ViT architecture [20, 39, 4, 14]; and long-context reasoning on Long Range Arena tasks (e.g., ListOps and Text Retrieval) [35]. Together, these benchmarks span bidirectional and autoregressive modeling, as well as moderate- and long-context text and image classification. During training, we treat $\beta$, introduced in Algorithm 1, as a trainable parameter. Beyond standard benchmarks, we conduct extreme-length scaling experiments on a single attention layer, reaching sequences of tens of millions of tokens while using the same hyperparameters as in the accuracy evaluations. Finally, FlashAttention-2/3 are exact fused implementations of Softmax Attention, and we use the terms interchangeably when reporting accuracy and runtime.
(a) GPU: RACE ($P=2,L=2$)
(b) GPU: RACE ($P=3,L=3$)
(c) GPU: RACE ($P=4,L=4$)
(d) GPU scaling
(e) CPU-RACE vs. GPU-Baselines
(f) CPU scaling
Figure 4: Scaling stress-tests on GPU and CPU: All plots use logarithmic axes and report a single forward–backward pass (batch size 1, 4 attention heads, d=128d{=}128) of a single attention layer. Linformer and Performer use the same hyperparameters as in Table 3. (a–d) compare the scaling behavior of RACE against attention baselines on GPU, while (e–f) show scaling for RACE (P=3,L=3P{=}3,L{=}3) on CPU, highlighting its superior efficiency at long sequence lengths.
Table 2: Comparison of attention methods across diverse tasks.
Method CIFAR-10 @1024 QNLI @2048 Tiny Stories @1024
Acc. (%) $\uparrow$ Acc. (%) $\uparrow$ PPL $\downarrow$
RACE (P=2, L=2) 63.7 60.7 4.2
RACE (P=3, L=3) 62.5 60.7 3.2
RACE (P=4, L=4) 65.9 61.1 2.6
Linear 60.0 60.7 7.0
Linformer-128 63.7 60.6 3.7
FlashAttention2 61.44 61.1 2.7
Angular ($\gamma$=8) 61.69 61.7 2.5
Performer-256 64.9 61.0 10.0
Sigmoid 57.2 61.1 3.7
Table 3: Food-101 @ 16K (Image Classification)
Method Train $\downarrow$ Test $\downarrow$ Acc. $\uparrow$
RACE (P=2, L=2) 891s 37s 42.4%
RACE (P=3, L=3) 950s 40s 43.5%
RACE (P=4, L=4) 1042s 42s 40.3%
Linear 1166s 44s 41.4%
Linformer-128 1250s 49s 20.2%
Performer-256 2546s 105s 42.4%
FlashAttention2 2600s 95s 42.1%
Baselines: We compare RACE against widely used baselines with publicly available implementations: FlashAttention2 [11], Linear Attention [18], Performer [7], Linformer [38], Sigmoid Attention [32], and YOSO [43]. These span exact, kernel-linear, and low-rank approximations. All models are tuned per authors’ guidelines and trained under identical settings.
Is RACE Attention as Accurate as Transformers? We report text-classification, image-classification, and masked language modeling results in Tables 1, 3, and 3.
Table 4: WikiText-103 @ 1024
Method PPL $\downarrow$
RACE (P=2, L=2) 23.9
RACE (P=2, L=3) 23.4
RACE (P=3, L=3) 21.9
RACE (P=3, L=4) 21.5
RACE (P=4, L=4) 20.9
FlashAttention2 20.9
Angular ($\gamma$=8) 19.0
Specifically, in Table 3, “Train/Test" denote per-epoch runtimes in seconds, and “Acc." denotes accuracy. Results for Linear Attention, Linformer, and Performer were collected using batch size 1 due to out-of-memory failures at batch size 8. In contrast, both RACE and FlashAttention2 remain memory-efficient at this sequence length and are evaluated with batch size 8. Additional experiments can be found in Appendix, refer to Tables 6, 7, 9, and 10. For autoregressive language modeling, RACE matches softmax-level perplexity on WikiText-103 and improves upon it on PTB (Tables 4 and 8). These results indicate that RACE preserves accuracy in the overlapping regime while delivering consistent gains on moderate- and long-context settings. Unless stated otherwise, all methods use the same Transformer backbone (layers, heads, embedding size, dropout) and training budget. We train with identical optimizers, schedulers, and batch sizes; full hyperparameters appear in Table 5. Metrics are reported from the best-validation checkpoint. All accuracy experiments were conducted on a single NVIDIA A100 GPU.
Can we reach 100 million context window on popular hardware? Now, we evaluate how RACE Attention scales across common hardware relative to strong baselines. For RACE, we use sketch parameters $(P,L)$ chosen to match FlashAttention2’s accuracy/perplexity on the same tasks in Table 3. For each method, we measure the wall-clock time for a single forward–backward pass of a single attention layer with 1 batch, 4 heads, and embedding size 128, as a function of sequence length, stress-testing context lengths up to 100 million tokens. Since FlashAttention-2/3 are designed specifically for GPU hardware, we use PyTorch’s optimized F.scaled_dot_product_attention implementation as the FlashAttention baseline when reporting CPU scaling results.
How far can we scale attention on standard Intel Xeon® Gold 5220R CPU? RACE scales to sequences of up to 75 million tokens for a single forward–backward pass of a single attention layer on CPU. By contrast, FlashAttention becomes prohibitively slow at $\sim\!2$ million tokens due to the quadratic time scaling in sequence length $N$ (see fig. 4(f)). It is worth noting that FlashAttention does not run out of memory on the CPU DRAM. RACE is more than $10000\times$ faster than FlashAttention at context length of $\sim\!33$ million. RACE finishes comfortably under 10 seconds for a single forward-backward pass on this hardware while FlashAttention takes approximately $10^{5}$ seconds on the same hardware. Although the absolute runtime grows with $N$, RACE’s speedup over FlashAttention increases. At 75 million tokens, RACE finishes in about 100 seconds. This is expected because RACE is linear and FlashAttention is quadratic in $N$. The experiments also highlight that linear attentions’ approximations are not only inaccurate but also significantly slower and have large memory overheads due to large hidden constants (see fig. 6 in Appendix). They run about an order of magnitude slower than RACE Attention and even go out of memory at $\sim\!33$ million tokens.
(a) GPU-RACE Speedup
(b) CPU-RACE Speedup
(c) YOSO vs. RACE on GPU
Figure 5: Speedup of RACE relative to GPU-only attention baselines, with RACE executed on GPU in (a) and on CPU in (b). (c) YOSO (CUDA kernel) fails beyond 32K tokens due to memory constraints, whereas RACE continues to scale to much longer sequences.
How far can we scale attention on the most powerful GH200 GPU? An NVIDIA GH200 has 96GB of memory. Here, we observe a similar trend. RACE scales up to 12 million tokens for a single forward-backward pass of a single attention layer, whereas FlashAttention-2/3 and Sigmoid Attention becomes impractical around $\sim\!4$ million tokens (see fig. 4). At $\sim\!4$ million tokens RACE takes merely 0.1 seconds to finish, while FlashAttention2 needs about 550 seconds, making RACE about $5500\times$ faster on GPUs for processing 4 million tokens. Additionally, RACE is about $5000\times$ faster than Sigmoid and $2600\times$ faster than FlashAttention3 for processing 4 million tokens (see fig. 5(a)). FlashAttention removes the quadratic memory cost of attention scores, but exact attention still requires storing token-level $Q,K,V,$ and $O$ activations and their gradients, incurring $\mathcal{O}(BHNd)$ memory. For large $N$, this linear footprint alone exceeds GPU HBM capacity. In contrast, RACE compresses queries and keys into $R$ bucket summaries per table, reducing activation memory to $\mathcal{O}(BHL(NR+Rd))$, which scales more favorably on GPU even than the linear attention baselines that exhaust memory at approximately $\sim 4$ million tokens (see fig. 4). As a result, RACE supports contexts up to $3.5\times$ longer than FlashAttention-2/3 and Sigmoid Attention.
Right Algorithm beats Hardware Acceleration: While GPUs offer substantial speedups for a fixed algorithm, comparing FlashAttention-2/3 and Sigmoid on a high-end GH200 GPU against RACE on a single CPU highlights the impact of algorithmic acceleration. Fig. 4(e) reports the runtime of a single forward–backward pass of a single attention layer as the context length increases. For short to moderate sequences ($N\lesssim 131\mathrm{K}$), the GPU’s massive parallelism dominates and FlashAttention-2/3 remain faster. Beyond this scale, however, the asymptotic behavior becomes decisive: the GPU kernels begin to saturate, and RACE becomes faster.At a sequence length of $\sim\!4$ million tokens (the largest supported by FlashAttention-2/3 in our configuration), RACE on CPU is roughly $40\times$ faster than FlashAttention-2 and Sigmoid Attention on GPU, and about $20\times$ faster than FlashAttention-3 (see fig. 5(b)). These results show that, in the long-context regime, state-of-the-art GPU attention kernels are fundamentally constrained by their quadratic dependence on sequence length. In contrast, RACE’s algorithmic efficiency enables it to outperform even the most advanced GPU-accelerated methods by a wide margin, despite executing on substantially weaker hardware.
@@ 5 Conclusion
We introduce RACE Attention, a linear-time, memory-efficient alternative to Softmax Attention that approximates a sharpened angular kernel via RACE sketches. This formulation will enable substantially more efficient key–value caching during inference and suggests a clear path toward optimized CUDA kernels that preserve favorable scaling behavior and theoretical guarantees in autoregressive settings. We leave a full exploration of these directions to future work.
@@ Acknowledgments
We are grateful to Aditya Desai and Zhaozhuo Xu for their valuable insights and thought-provoking discussions. The work was partially supported by Rice Ken Kennedy Institute (K2I) Generative AI Cluster Funding.
@@ 5.1 Additional Notes on Experiments
Unless otherwise specified, we adopt a linear warmup–decay learning-rate schedule. Let $T$ denote the total number of optimizer updates,
$T=\text{epochs}\times\text{len(train\_loader)}.$
The learning rate increases linearly from 0 to the base learning rate (reported in Table 5) over the first $0.01T$ updates, and then decays linearly to 0 over the remaining T0.01TT-0.01T updates. The learning-rate scheduler is stepped once per optimizer update. We use the AdamW optimizer with hyperparameters listed in Table 5. For experiments with a learning rate below $2\times 10^{-4}$, we do not apply linear warmup or a learning-rate scheduler; instead, we train the model using a constant learning rate for the number of epochs specified in Table 5.
Data Pre-processing for Arxiv: To evaluate long-context classification performance, we construct 16K-, 32K-, and 64K-token variants of the ArXiv dataset using a two-stage preprocessing pipeline. First, we tokenize all documents with a basic English tokenizer and retain only those whose raw token length exceeds a predefined minimum threshold, ensuring that each example contains sufficiently long context. Next, we perform streaming sequence packing. Documents are grouped by class and concatenated sequentially until a target sequence length of 16K, 32K, or 64K tokens is reached. To maintain high packing efficiency, only very small residual fragments are discarded. This procedure yields long, contiguous sequences that preserve the natural structure of individual documents while enabling fixed-length training. After balancing classes, the resulting dataset sizes are as follows: 16K (1,947 training examples, 209 test examples), 32K (3,650 training examples, 520 test examples), and 64K (3,882 training examples, 384 test examples).
Data Pre-processing for Food-101:
To evaluate long-context image classification on the Food-101 dataset with a 16K token context length, we reduce the dataset size by discarding 50 classes. We train the model on 20,000 samples and evaluate it on a held-out test set of 2,500 samples. Images are processed using a patch size of 4, an input resolution of $512\times 512$, and a stride of 4. All remaining hyperparameters are reported in Table 5.
Table 6: FashionMNIST @ 784
Method Accuracy
RACE (P=2, L=5) 87.7%
RACE (P=3, L=5) 87.5%
RACE (P=4, L=4) 86.6%
RACE (P=4, L=5) 85.7%
Linformer-128 87.7%
FlashAttention2 87.2%
Angular ($\gamma$=8) 86.4%
Linear 85.8%
Performer-256 86.6%
Table 7: Tiny Stories @ 512
Method Perplexity
RACE (P=3, L=4) 3.9
RACE (P=4, L=4) 3.3
RACE (P=5, L=4) 2.7
RACE (P=5, L=5) 5.1
Linear 6.0
Angular ($\gamma$=8) 2.9
FlashAttention2 3.1
Linformer-128 4.6
Performer-256 7.1
Table 8: Penn Tree Bank @ 128
Method Test PPL
RACE (P=2, L=2) 54.7
RACE (P=3, L=3) 54.2
RACE (P=4, L=4) 53.4
Angular ($\gamma$=8) 58.8
Angular ($\gamma$=12) 57.6
Linear 73.2
FlashAttention2 55.4
Table 9: Comparison of attention methods on Long Range Arena
Method ListOps @2000 Text Retrieval @8000
Acc. (%) $\uparrow$ Acc. (%) $\uparrow$
RACE (P=2, L=2) 41.9% 80.3%
RACE (P=3, L=3) 41.3% 80.8%
RACE (P=4, L=4) 41.6% 80.9%
Linformer-128 38.9% 76.1%
FlashAttention2 41.4% 80.5%
Angular ($\gamma$=8) 42.2% OOM
Linear 39.6% 80.6%
Performer-256 40.2% 80.8%
YOSO-16 41.0% 80.2%
Table 10: Comparison of attention methods on natural language understanding tasks
Method Yahoo @256 IMDB @512 SST-2 @1024
Acc. (%) $\uparrow$ Acc. (%) $\uparrow$ Acc. (%) $\uparrow$
RACE (P=2, L=2) 66.9% 80.6% 76.7%
RACE (P=3, L=3) 66.6% 81.3% 77.1%
RACE (P=4, L=4) 67.2% 81.3% 79.4%
Linformer-128 64.7% 78.2% 75.1%
FlashAttention2 67.2% 80.0% 78.5%
Angular ($\gamma$=8) 67.0% 79.6% 77.2%
Linear 66.9% 80.9% 78.0%
Performer-256 64.9% 81.0% 77.3%
(a) CPU: RACE ($P=2,L=2$)
(b) CPU: RACE ($P=3,L=3$)
(c) CPU: RACE ($P=4,L=4$)
(d) $d$ scaling @ N=65536N{=}65536
(e) GPU: RACE ($P=4,L=4$)
Figure 6: (a–c) CPU results for different $(P,L)$. (d) Linear attention scales quadratically in $d$, while RACE scales linearly (fixed N=65536N{=}65536). (e) GPU scaling under hyperparameters different from fig. 4.